HIGH RESOLUTION INFRARED DIODE LASER SPECTROSCOPY OF X^{2}Π CBr

\underline{ANDREW J. MARR} and TREVOR J. SEARS, Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973-5000; PAUL B. DAVIES, University of Cambridge, Lensfield Road, Cambridge, United Kingdom CB2 1EW.

The infrared spectrum of the X^{2}Π state of CBr from 680 - 750 cm^{−1} has been measured using diode laser absorption spectroscopy. The CBr radicals were produced by 193 nm excimer photolysis of bromoform. Transitions belonging to the v = 1 ← 0, 2 ← 1, 3 ← 2 and 4 ← 3 bands of both isotopomers, C^{79}Br and C^{81}Br have been identified. Employing case(a) matrix elements, spectroscopic parameters were obtained from a least squares fit where the transition frequencies of both isotopomers were fitted simultaneously. The results were used to obtain a Rydberg-Klein-Rees (RKR) potential. Additionally, the kinetics of the CBr + O_{2} reaction were investigated. This is the first recording of vibrational spectra of CBr.

Acknowledgements: This work was carried out at Brookhaven National Laboratory under Contract No. DE-AC02-76CH00016 with the U.S. Department of Energy and support by its Division of Chemical Sciences, Office of Basic Energy Sciences. Support was also provided by a NATO Collaborative Research Grant, No. CRG 940066.